Gdub08:

If 4x + 5 = 13x + 4 – x – 9, then x = ?

Gdub08:

Are you supposed to combine like terms? the 4x and the 13x?

umm:

Yes, combine both terms. But firstly, you need to simplify both sides of the equation, like so \[4 x + 5 = 13 x + 4 − x − 9\] \[4 x + 5 = 13 x + 4 + − x + − 9\] Here you combine terms \[4 x + 5 = ( 13 x + − x ) + ( 4 + − 9 )\]

Gdub08:

@umm wrote:
Yes, combine both terms. But firstly, you need to simplify both sides of the equation, like so \[4 x + 5 = 13 x + 4 − x − 9\] \[4 x + 5 = 13 x + 4 + − x + − 9\] Here you combine terms \[4 x + 5 = ( 13 x + − x ) + ( 4 + − 9 )\]
so then x = 5/4, right?... I think. or 1 1/4